OpenStudy (anonymous):

Find all solutions to the equation in the interval [0, 2π) sin 2x = -sin x

OpenStudy (anonymous):

sin 2x = 2sin x cos x

OpenStudy (anonymous):

from the prob, sin 2x=-sin x, rearranging the term.... sin 2x + sin x =0

OpenStudy (anonymous):

now apply the identity... 2 sin x cos x + sin x =0 sin x(2cos x + 1) =0 , hence sin x=0 and 2 cos x +1 =0 x=arc sin 0 and 2cosx=-1, cos x= (-1/2) x= arc cos(-1/2) x=0 and x=2pi/3.., where pi=3.1416